### v1

# CALT-TH-2024-013

Abstract
Motivated by string-theoretic swampland conjectures, the existence of a dark fifth dimension, whose size is roughly 1 – 10 microns, has been proposed. A great deal of supporting evidence has been presented, and definitive experimental tests are likely to be carried out. The basic idea is that the four-dimensional spacetime that we observe lives on a brane that is localized in the dark dimension. This short note points out that there are two distinct ways to realize such a scenario in string theory/M-theory. In the one considered previously the dark dimension is topologically a circle and our observable 4d spacetime is confined to a brane that is localized in a GUT-scale region of the circle. An alternative possibility is that the dark dimension is a line interval with branes attached at each end. This option would imply the existence of a parallel 4d spacetime microns away from us!
Motivated by string-theoretic swampland conjectures [1], it has recently been proposed that there is a mesoscopic fifth dimension whose size is roughly 1 – 10 microns [2][6]. This dimension is called dark, because it only supports gravitational forces. The basic idea is that the four-dimensional spacetime that we observe lives on a brane that is localized at a point (more precisely, a GUT-scale region) in the dark dimension (DD). Any additional compact dimensions must be very much smaller than the DD. Thus, for many purposes, we may assume that the brane is well approximated by a standard-model 3-brane (SM3-brane) even though it would be desirable to understand exactly how such a brane arises from a more fundamental string theory/M-theory starting point. For reasons explained in [2], the existence of more than one mesoscopic dimension is excluded. As discussed in [3], astrophysical constraints limit the allowed size of the DD to 1 – 10 microns, and future large-scale structure surveys could support or rule out its existence.
A number of arguments were made in [2][6], based on various theoretical and observational facts, in support of the DD proposal. These arguments suggest that the size of the DD is not very far below the limit (about 30 microns) obtained by a recent experimental study of Newton’s $1/r^{2}$ gravitational force law at very short distances [7]. That experiment used “a stationary torsion-balance detector and a rotating attractor containing test bodies.” An experiment that can probe even shorter distances is being planned [8][9]. It will study the orbits of a microscopic planetary system. Precession of the orbit would imply a central potential different from $1/r$. Other experimental approaches may be possible, e.g. based on atom interferometry.
It is now conceivable that the DD realization of string theory will receive experimental support within this decade. Prior to the DD proposal I did not think that string theory could make novel testable predictions in the foreseeable future. This DD proposal appears to be well motivated, but even so it should be considered speculative. The experiments ought to be a very high priority, because experimental verification would be revolutionary. A negative result would not be revolutionary, but it would still be fundamental and very important.
There have been previous theoretical proposals for a relatively large extra dimension, such as [10] and [11]. They differ in important respects from the current proposal, which is motivated by swampland considerations as well as various astrophysical arguments. In particular, the current proposal shows promise for giving a deeper quantitative understanding of dark matter and dark energy without invoking the anthropic principle.
If a mesoscopic DD is discovered, it will become important to figure out how it fits into string theory. We have known since the mid-1990s that the various versions of superstring theory/M-theory are related by dualities, so that there really is a unique underlying theory. However, the different versions, which focus on different corners of the moduli space of vacua, can appear more or less promising as starting points for phenomenology. There have been two such classes of starting points that account for almost all studies of string phenomenology. One class is based on the $E_{8}\times E_{8}$ heterotic string [12]. The other class is based on F-theory [13], which utilizes certain nonperturbative solutions of type IIB superstring theory. See [14] for a recent discussion of the status of string phenomenology.
An F-theory approach to realizing the DD proposal is described in [6]. In that realization the DD is topologically a circle, and the SM3-brane is localized in a GUT-scale region of the circle. That paper argues that this picture gives a tower of states, associated to a 7-brane, at $10^{9}$$10^{10}$ GeV. This scale, called the species scale, corresponds to the 5d Planck scale. There should also be a GUT scale a few orders of magnitude higher. The SM3-brane approximation should be sufficient for phenomenology in the foreseeable future, though a deeper understanding of how the SM3-brane arises would be desirable.
Following the discovery that cancellation of local gauge and gravitational anomalies in 10d superstring theories is only possible for the gauge groups $SO(32)$ and $E_{8}\times E_{8}$ [15], a string theory with the latter gauge group, called the $E_{8}\times E_{8}$ heterotic string, was constructed [12]. For the following decade (prior to the introduction of F-theory) this was the only version of superstring theory that showed promise for giving realistic phenomenology. The introduction of Calabi–Yau compactification [16] launched a great deal of interest in superstring theory in general, and this specific approach in particular. It was clear from the outset that this setup has many qualitative successes. However, even today there is no specific construction that looks exactly right. The math is hard, and the number of possibilities is very large. The same can be said for the F-theory approach.
Among the many exciting discoveries made in the mid-1990s was the realization that the $E_{8}\times E_{8}$ heterotic string theory and the type IIA superstring theory actually have an eleventh dimension [17][18] that is not visible in perturbation theory. In the type IIA case there is a circular eleventh dimension ($S^{1}$) whose size is proportional to the string coupling constant. In the limit of infinite coupling one obtains 11-dimensional M-theory. In the case of the $E_{8}\times E_{8}$ theory there is a semi-circular eleventh dimension ($S^{1}/Z_{2}$), whose size is proportional to the string coupling constant. This dimension can also be viewed as a line interval with two ends. Each of the ends is attached to an end-of-the-world 9-brane (an EW9-brane). One $E_{8}$ gauge field is associated to each of the two EW9-branes. This is required by the anomaly analysis [17]. The 11-dimensional bulk, in between the EW9-branes, is described by 11-dimensional M-theory.
The eleventh dimension of the $E_{8}\times E_{8}$ theory is a DD. Most studies of Calabi–Yau compactification of this theory have ignored the eleventh dimension, which is a reasonable thing to do for many purposes. However, it is an important fact that this class of string theory solutions necessarily implies the existence of a DD. What is new now, is the suggestion that this dimension should be several microns in length.
Suppose we consider implementing the DD proposal using the $E_{8}\times E_{8}$ theory. Then an essential difference from the F-theory example is that it requires the existence of two EW-branes. After a suitable compactification of six dimensions they can be approximated by 3-branes, and one of them can be identified as the SM3-brane that describes our 4d reality at least up to energies of a TeV and possibly much further. In this setup there is a second parallel 3-brane, just microns away! Should we expect the physics of the second EW-brane to be the same as or different from “our” EW-brane? Much of the Calabi–Yau phenomenology has found it useful to implement supersymmetry breaking in the invisible $E_{8}$, by gaugino condensation for example, and for this breaking to be transferred to the other observable $E_{8}$ factor gravitationally. If this is correct, it would imply that the physics of the two EW-branes is different. Whether or not that is the case, the possibility of a whole other parallel 4d spacetime just microns away from us is challenging to visualise and seems quite incredible! This is not a prediction of the DD proposal – just a possible option.
For strong coupling the $E_{8}\times E_{8}$ heterotic string theory is not a string theory. (A similar remark can be made for the type IIA superstring theory.) In M theory the basic branes are an M2-brane and its magnetic dual, which is an M5-brane. A cylindrical M2-brane can be stretched across the dark dimension with its boundaries attached to the EW-branes. In the weak-coupling limit the two EW-branes approach one another, and the cylinder collapses to become the $E_{8}\times E_{8}$ string. However, if the branes are microns apart, it looks nothing like a string – unless it can be approximated by an open string stretched between the two EW-branes. In the $E_{8}\times E_{8}$ theory the coupling constant $g$ is proportional to the size of the dark dimension, $l_{d}$. More specifically, $g\sim l_{d}/l_{s}$, where $l_{s}$ is a string or Planck scale, which is very small. If $l_{d}$ is several microns, this ratio is enormous. The problem is evident from the shape of the stretched M2-brane. If there were an S-dual formulation of the $E_{8}\times E_{8}$ heterotic string theory, whose coupling constant is $\tilde{g}=1/g$, that could be helpful. The $SO(32)$ heterotic string theory has an S-dual formulation, namely the type I superstring theory, which is based on a non-BPS (breakable) string. It seems unlikely that there is an analogous S-dual of the $E_{8}\times E_{8}$ heterotic string theory. The problem is that it seems impossible to define a $g\to\infty$ limit while retaining the EW-branes.
Suppose that experimental evidence for a DD is obtained. How could we decide whether the DD is a circle or a line segment with a second brane at the other end of the DD? We would only be sensitive to the gravitational effects of the other brane. One could imagine a star on our brane orbiting an invisible object that resides on the other brane. Would it be possible to establish that the invisible object is not a black hole or something on our brane that we already understand? Alternatively, it is conceivable that a very sensitive test of gravity at short distances could distinguish between the two alternatives.
The original papers [2][6] make a strong case for Kaluza–Klein excitations of gravitons in the DD comprising the most important component of dark matter. Of course, it is still possible that there are also other sources of dark matter. Additional sources of dark matter in the DD might include KK excitations of sterile neutrinos. Possible sources of dark matter on the SM3-brane are the familiar ones that have received the most attention by both theorists and experimentalists so far. These include axions and wimps, though the latter have been largely excluded. If there is a second EW-brane, matter that is confined to that brane would be dark matter from our point of view.
All known black holes are much larger than the proposed DD. They should appear the same to observers in either of the two EW-branes (if there are two). Black holes that are smaller than the DD would be localized within that dimension, away from any branes, and they would be very difficult to detect. The implications of the DD proposal for primordial black holes have been discussed in [19]-[21]. These papers point out that in the DD scenario 5d black holes have a much longer lifetime due to Hawking decay than 4d ones of the same mass. Therefore, they suggest that primordial black holes that are smaller than the DD could be another significant component of dark matter.
In conclusion, the DD scenario is a well-motivated proposal, based on plausible string theory considerations, that should be experimentally testable. Its confirmation would be revolutionary for particle physics and astrophysics. If that happens, an important question would be whether or not there is a parallel brane microns away from our SM3-brane.
Acknowledgments
I am grateful to Cumrun Vafa for encouraging me to release this note. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of High Energy Physics, under Award Number DE-SC0011632.